# HAPNet: Toward Superior RGB-Thermal Scene Parsing via Hybrid, Asymmetric, and Progressive Heterogeneous Feature Fusion

### v2

@@ III-B2 Cross-modal Spatial Prior Descriptor
Compared to Transformers [41, 10], CNNs have demonstrated superior performance in capturing local semantics, underscoring their significance in computer vision applications that necessitate clear object boundaries. ConvNeXt [14] have shown exceptional performance in capturing rich, robust visual features in comparison with other hierarchical CNN architectures such as ResNet [12] and its variants [42], as evidenced by recent endeavors using ConvNeXt for multi-modal scene parsing [43, 44]. Given these advantages, we employ basic ConvNeXt blocks to construct our CSPD. Furthermore, incorporating RGB data as a complementary input enables the CSPD to extract more comprehensive local spatial patterns than using thermal images alone. Consequently, we construct our CSPD using two identical, weight-sharing subnetworks, enabling efficient and effective extraction of cross-modal spatial priors from the RGB-T data.
@@
Specifically, $\boldsymbol{I}^{R}$ and $\boldsymbol{I}^{T}$ are independently fed into a series of weight-sharing ConvNeXt blocks, generating multi-scale features $\mathcal{P}^{R}=\{\boldsymbol{P}^{R}_{2},\boldsymbol{P}^{R}_{3},\boldsymbol{P}^{R}_{4}$} and $\mathcal{P}^{T}=\{\boldsymbol{P}^{T}_{2},\boldsymbol{P}^{T}_{3},\boldsymbol{P}^{T}_{4}$}, respectively, where $\boldsymbol{P}^{R,T}_{i}\in\mathbb{R}^{\frac{H}{S_{i}}\times\frac{W}{S_{i}}\times C_{i}}$ represents the features at $i$-th stage, $C_{i}$ and $S_{i}=2^{i+1}$ ($i\in[2,4]\cap\mathbb{Z}$) denote the corresponding channel and stride numbers, respectively. We then combine $\boldsymbol{P}^{R}_{i}$ and $\boldsymbol{P}^{T}_{i}$ via element-wise summation, and reduce the results to $D$ channels through $1\times 1$ convolutions, yielding heterogeneous features $\mathcal{P}^{H}=\{\boldsymbol{P}^{H}_{2},\boldsymbol{P}^{H}_{3},\boldsymbol{P}^{H}_{4}$}, where $\boldsymbol{P}^{H}_{i}\in\mathbb{R}^{\frac{H}{S_{i}}\times\frac{W}{S_{i}}\times D}$. The three-scale features are then flattened and concatenated to form the cross-modal spatial prior $\boldsymbol{F}^{P}_{1}$ which is used for the following encoding stages. We present detailed ablation studies in Sect. IV-E to analyze the impact of different CSPD construction blocks and data input strategies on the overall performance.
III-B3 Progressive Heterogeneous Feature Integrator
@@
where $\mathrm{MHA}(\cdot)$ represents the multi-head attention (MHA) operation (implemented based on multi-scale deformable attention [36] to reduce computations) and $\mathrm{LN}(\cdot)$ represents the layer normalization operation. Following common practices in attention modules such as [45, chen2022vision, 43], we introduce a learnable coefficient $\kappa_{i}$ to dynamically adjust the weight of the MHA, enabling flexible fusion of the local semantics-enhanced features into the VFM. After the $i$-th ViT block, the output feature maps $\boldsymbol{F}^{V}_{i+1}$ with enriched global context are yielded. $\boldsymbol{F}^{V}_{i+1}$ is then fed into the CCG to fuse its global context with the spatial prior $\boldsymbol{F}^{P}_{i}$. In this process, $\boldsymbol{F}^{P}_{i}$ serves as the query, while $\boldsymbol{F}^{V}_{i+1}$ acts as the keys and values to perform MHA within the CCG as follows:
@@
After the MHA operation, a feed-forward network (FFN) is applied to further process the fused features, resulting in the updated spatial prior $\boldsymbol{F}^{P}_{i+1}$, which have already incorporated rich local and global semantics and will be inputted to the next encoding stage.
III-C Mask Classification Decoder
@@
III-D Auxiliary Local Semantics Enhancement Task
In our mask classification decoder, $\boldsymbol{E}^{P}$ generated from $\boldsymbol{F}^{F}_{\frac{1}{4}}$ typically provides rich per-pixel semantics of the scene, while $\boldsymbol{E}^{M}$ generated from the other three scales of features provides information on category-related cluster centers [46]. To ensure that $\boldsymbol{E}^{P}$, utilized by our mask classification decoder, contains robust and distinguishable local semantics, we incorporate deep supervision into $\boldsymbol{F}^{F}_{\frac{1}{4}}$. Specifically, we attach $\boldsymbol{F}^{F}_{\frac{1}{4}}$ with a lightweight fully convolutional decoder head and directly supervise this auxiliary network to produce semantic predictions. After two convolutional layers, the auxiliary semantic predictions with $(N-1)$ channels and the same resolution as $\boldsymbol{F}^{F}_{\frac{1}{4}}$ are generated (excluding the “no object” category). This design enables the corresponding encoder layers to produce a $\boldsymbol{F}^{F}_{\frac{1}{4}}$ with more distinguishable local semantics, further ensuring that the pixel decoder can generate a more distinguishable $\boldsymbol{E}^{P}$, and ultimately leading to improved scene parsing performance without increasing network parameters or computational complexity.
@@ III-D Auxiliary Local Semantics Enhancement Task
III-E Loss Function
We use a combination of the cross-entropy (CE) loss $\mathcal{L}_{\text{cls}}$ for class prediction, along with the binary CE loss $\mathcal{L}_{\text{bce}}$ and the dice loss $\mathcal{L}_{\text{dice}}$ for mask prediction. For our local semantics enhancement task, we directly calculate the standard CE loss $\mathcal{L}_{\text{ce}}$ over its semantic predictions. Therefore, the overall loss function can be expressed as follows:
$\mathcal{L}=\lambda_{\text{bce}}\mathcal{L}_{\text{bce}}+\lambda_{\text{dice}}\mathcal{L}_{\text{dice}}+\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}+\lambda_{\text{ce}}\mathcal{L}_{\text{ce}}.$ (3)
Following [36], we set $\lambda_{\text{bce}}=5.0$, $\lambda_{\text{dice}}=5.0$, $\lambda_{\text{cls}}=2.0$, and $\lambda_{\text{ce}}=0.4$, respectively, and we set $\lambda_{\text{cls}}=0.1$ for the “no object” category. Additionally, we employ bipartite matching [47] to determine the least-cost assignment during model training.
TABLE I: Quantitative comparisons (%) with the SoTA RGB-T scene parsing methods on the MFNet test set. The symbol ‘-’ denotes missing data in the original publication, and the best results are presented in bold font.
@@ IV Experiments
IV-A Datasets and Experimental Setup
@@
TABLE II: Quantitative comparisons (%) with the SoTA RGB-T scene parsing methods on the KP day-night test set. The best results are presented in bold font.
@@ Methods
IV-A1 MFNet [5]
IV-A2 PST900 [1]
TABLE III: Quantitative comparisons (%) with the SoTA RGB-T scene parsing methods on the PST900 test set. The symbol “-” denotes missing data in the original publication, and the best results are presented in bold font.
Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU
MFNet [5] - 98.6 - 41.1 - 64.2 - 60.3 - 20.7 - 57.0
RTFNet [4] - 98.9 - 36.4 - 75.3 - 52.0 - 25.3 - 57.6
EGFNet [48] - 99.2 - 74.3 - 83.0 - 71.2 - 64.6 - 78.5
ABMDRNet [6] - 98.7 - 24.1 - 72.9 - 54.9 - 57.6 - 67.3
FEANet [21] - - - - - - - - - - 91.4 85.5
DBCNet [2] - 98.9 - 62.3 - 71.1 - 52.4 - 40.6 - 74.5
CAINet [10] 99.6 99.5 95.8 80.3 96.0 88.0 88.3 77.2 91.3 78.6 94.2 84.7
EAEFNet [3] 99.8 99.5 92.2 80.4 91.0 87.7 93.0 83.9 79.3 75.6 91.1 85.4
GMNet [9] 99.8 99.4 90.2 85.1 89.0 83.8 88.2 73.7 80.8 78.3 89.6 84.1
CRM-RGBTSeg [7] - 99.6 - 79.5 - 89.6 - 89.0 - 82.2 - 88.0
TABLE IV: Quantitative comparisons (%) with the SoTA data-fusion scene parsing networks on the NYU-Depth V2 test set.
@@ IV-A2 PST900 [1]
TABLE V: An ablation study (%) on different data input strategies on the MFNet test set.
@@
TABLE VI: An ablation study (%) on different CSPD construction blocks on the MFNet test set.
@@
Duplex ResNet sharing weights 66.3 48.0 38.4 19.9 9.1 0.0 29.6 23.6 42.3 36.8
Duplex Swin-Transformer sharing weights 89.7 74.1 67.2 47.6 31.0 6.3 52.7 57.0 74.6 58.2
Duplex ConvNeXt with separate weights 89.6 74.6 66.6 46.3 45.7 10.2 56.6 62.5 73.6 61.2
Duplex ConvNeXt sharing weights (Ours) 89.1 76.0 67.2 50.6 41.3 7.9 58.6 58.7 75.1 60.9
TABLE VII: An ablation study (%) on the effectiveness of GLCA and CCG on the MFNet test set. When both components are removed, the spatial priors extracted using CSPD and the global context extracted using ViT are fused via element-wise summation after resolution alignment.
@@
TABLE VIII: An ablation study (%) on different duplex encoder architectures on the MFNet test set.
@@
Duplex BEiTv2 concatenation 65.7 57.4 12.7 27.7 8.5 0.0 14.7 20.6 40.9 33.7
Duplex BEiTv2 summation 65.4 57.1 23.3 28.0 8.5 0.0 19.9 16.2 42.6 35.0
Duplex ConvNeXt concatenation 89.7 74.1 66.4 45.3 29.8 4.8 52.8 55.5 69.2 57.4
Duplex ConvNeXt summation 90.7 74.6 66.9 46.8 32.9 6.6 56.6 52.7 72.0 58.5
@@
TABLE IX: An ablation study (%) on different VFMs on the MFNet test set. “MM” represents multi-modal pre-training. BEiT and BEiTv2 are trained via a self-supervised learning strategy named masked image modeling, while DINOv2 is trained via discriminative self-supervised learning strategy.
@@
DeiT [55] Supervised ImageNet-1K 90.4 75.2 66.6 47.6 36.3 0.7 59.9 60.0 68.3 59.4
AugReg [56] Supervised ImageNet-22K 90.3 74.9 66.8 48.2 35.8 6.0 56.6 56.5 74.7 59.3
BEiT [16] Self-Supervised ImageNet-22K 90.7 75.0 67.6 48.6 39.1 4.2 55.0 59.0 72.5 59.7
DINOv2 [18] Self-Supervised LVD-142M (MM) [18] 90.0 74.7 66.8 45.1 42.1 9.7 56.2 59.7 76.4 60.3
BEiTv2 [17] Self-Supervised ImageNet-22K (MM) 89.1 76.0 67.2 50.6 41.3 7.9 58.6 58.7 75.1 60.9
@@ IV-A4 NYU-Depth V2 [59]
This dataset has been widely used for the evaluation of indoor scene parsing performance evaluation. It provides $1,449$ pairs of RGB and depth images (resolution: $480\times 640$ pixels), and semantic annotations for $40$ categories. We adopt the same dataset splitting strategy as presented in [40].
IV-B Evaluation Metrics

### v3

@@ 3.2.2 Cross-modal Spatial Prior Descriptor
Compared to Transformers [52, 10], CNNs have demonstrated superior performance in capturing local semantics, underscoring their significance in computer vision applications that necessitate clear object boundaries. ConvNeXt [17] has shown exceptional performance in capturing rich, robust visual features in comparison with other hierarchical CNN architectures such as ResNet [14] and its variants [53], as evidenced by recent endeavors using ConvNeXt for multi-modal scene parsing [54, 55]. Given these advantages, we employ basic ConvNeXt blocks to construct our CSPD. Furthermore, incorporating RGB data as a complementary input enables the CSPD to extract more comprehensive local spatial patterns than using thermal images alone. Consequently, we construct our CSPD using two identical, weight-sharing subnetworks, enabling efficient and effective extraction of cross-modal spatial priors from the RGB-T data.
@@
Specifically, $\boldsymbol{I}^{R}$ and $\boldsymbol{I}^{T}$ are independently fed into a series of weight-sharing ConvNeXt blocks, generating multi-scale features $\mathcal{P}^{R}=\{\boldsymbol{P}^{R}_{2},\boldsymbol{P}^{R}_{3},\boldsymbol{P}^{R}_{4}$} and $\mathcal{P}^{T}=\{\boldsymbol{P}^{T}_{2},\boldsymbol{P}^{T}_{3},\boldsymbol{P}^{T}_{4}$}, respectively, where $\boldsymbol{P}^{R,T}_{i}\in\mathbb{R}^{\frac{H}{S_{i}}\times\frac{W}{S_{i}}\times C_{i}}$ represents the features at $i$-th stage, $C_{i}$ and $S_{i}=2^{i+1}$ ($i\in[2,4]\cap\mathbb{Z}$) denote the corresponding channel and stride numbers, respectively. We then combine $\boldsymbol{P}^{R}_{i}$ and $\boldsymbol{P}^{T}_{i}$ via element-wise summation, and reduce the results to $D$ channels through $1\times 1$ convolutions, yielding heterogeneous features $\mathcal{P}^{H}=\{\boldsymbol{P}^{H}_{2},\boldsymbol{P}^{H}_{3},\boldsymbol{P}^{H}_{4}$}, where $\boldsymbol{P}^{H}_{i}\in\mathbb{R}^{\frac{H}{S_{i}}\times\frac{W}{S_{i}}\times D}$. The three-scale features are then flattened and concatenated to form the cross-modal spatial prior $\boldsymbol{F}^{P}_{1}$ which is used for the following encoding stages. We present detailed ablation studies in Sect. 4.5 to analyze the impact of different CSPD construction blocks and data input strategies on the overall performance.
3.2.3 Progressive Heterogeneous Feature Integrator
@@
where $\mathrm{LN}(\cdot)$ represents the layer normalization operation, $\mathrm{MHA}(\cdot,\cdot)$ represents the multi-head attention operation. Following common practices in attention modules such as [56, 57, 54], we introduce a learnable coefficient $\kappa_{i}$ to dynamically adjust the weight of the MHA, enabling flexible fusion of the local semantics-enhanced features into the VFM. After the $i$-th ViT block, the output feature maps $\boldsymbol{F}^{V}_{i+1}$ with enriched global context are yielded. $\boldsymbol{F}^{V}_{i+1}$ is then fed into the CCG to fuse its global context with the spatial prior $\boldsymbol{F}^{P}_{i}$. In this process, $\boldsymbol{F}^{P}_{i}$ serves as the query, while $\boldsymbol{F}^{V}_{i+1}$ acts as the keys and values to perform MHA within the CCG as follows:
@@
After the MHA operation, a feed-forward network (FFN) is applied to further process the fused features, resulting in the updated spatial prior $\boldsymbol{F}^{P}_{i+1}$, which has already incorporated rich local and global semantics and will be inputted to the next encoding stage. In the above two components, the MHA processes are implemented based on multi-scale deformable attention (MSDA) [31] to reduce computations. Please refer to the supplemental material for more detailed mathematical derivations of MSDA.
3.3 Mask Classification Decoder
@@
3.4 Auxiliary Local Semantics Enhancement Task
In our mask classification decoder, $\boldsymbol{E}^{P}$ generated from $\boldsymbol{F}^{F}_{\frac{1}{4}}$ typically provides rich local (per-pixel) semantics of the scene, while $\boldsymbol{E}^{M}$ generated from the other three scales of features provides information on category-related cluster centers [58]. To enhance the robustness and discriminability of local semantics utilized by our mask classification decoder, we incorporate deep supervision into $\boldsymbol{F}^{F}_{\frac{1}{4}}$, effectively reducing noisy and ambiguous patterns within it. Specifically, we attach $\boldsymbol{F}^{F}_{\frac{1}{4}}$ with a lightweight fully convolutional decoder head and directly supervise this auxiliary network to produce semantic predictions. After two convolutional layers, the auxiliary semantic predictions with (N1)(N-1) channels and the same resolution as $\boldsymbol{F}^{F}_{\frac{1}{4}}$ are generated (excluding the “no object” category). This design enables the corresponding encoder layers to produce a $\boldsymbol{F}^{F}_{\frac{1}{4}}$ with more distinguishable local semantics, further ensuring that the pixel decoder can generate a more distinguishable $\boldsymbol{E}^{P}$, and ultimately leading to improved scene parsing performance without increasing network parameters or computational complexity.
@@ 3.4 Auxiliary Local Semantics Enhancement Task
3.5 Loss Function
We employ the cross-entropy (CE) loss $\mathcal{L}_{\text{ce}}$ for the class prediction of object queries, along with the binary cross-entropy (BCE) loss $\mathcal{L}_{\text{bce}}$ and dice loss $\mathcal{L}_{\text{dice}}$ for mask prediction. $\mathcal{L}_{\text{ce}}$ can be formulated as follows:
$\mathcal{L}_{\text{ce}}=-\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{y}_{c}^{i}\log(\boldsymbol{p}_{c}^{i}),$ (3)
where $N$ denotes the total number of object queries, $\boldsymbol{y}_{c}^{i}$ denotes the one-hot encoded ground-truth vector for $c$ categories, and $\boldsymbol{p}_{c}^{i}$ stores values indicating the prediction probabilities over $c$ categories. $\mathcal{L}_{\text{bce}}$ and $\mathcal{L}_{\text{dice}}$ for mask classification are formulated as follows:
$\mathcal{L}_{\text{bce}}=-\frac{1}{M}\sum_{i=1}^{M}(y_{b}^{i}\log(p_{b}^{i})+(1-y_{b}^{i})\log(1-p_{b}^{i})),$ (4)
$\mathcal{L}_{\text{dice}}=1-\frac{2\sum_{i=1}^{M}y_{b}^{i}p_{b}^{i}}{\sum_{i=1}^{M}y_{b}^{i}+\sum_{i=1}^{M}p_{b}^{i}},$ (5)
where $M$ denotes the total number of pixels, $y_{b}^{i}\in\{0,1\}$ denotes the binary ground truth of one foreground pixel, and $p_{b}^{i}\in[0,1]$ is a scalar indicating the corresponding prediction probability. For our local semantics enhancement task, we directly calculate another CE loss $\mathcal{L}_{\text{aux}}$ over its semantic predictions. The overall loss function can be expressed as follows:
$\mathcal{L}=\lambda_{\text{bce}}\mathcal{L}_{\text{bce}}+\lambda_{\text{dice}}\mathcal{L}_{\text{dice}}+\lambda_{\text{ce}}\mathcal{L}_{\text{ce}}+\lambda_{\text{aux}}\mathcal{L}_{\text{aux}}.$ (6)
We follow [45] and [37] to set $\lambda_{\text{bce}}=5.0$, $\lambda_{\text{dice}}=5.0$, $\lambda_{\text{ce}}=2.0$, and $\lambda_{\text{aux}}=0.4$, respectively, and we set $\lambda_{\text{ce}}=0.1$ for the “no object” category. Additionally, we employ bipartite matching [59] to determine the least-cost assignment during model training.
Table 1: Quantitative comparisons (%) with the SoTA RGB-T scene parsing methods on the MFNet test set. The symbol ‘-’ denotes missing data in the original publication, and the best results are presented in bold font. The values of Acc and IoU for the “background” category are omitted in the table, but they are still included in the calculation of the corresponding mean values.
@@ 4 Experiments
4.1 Datasets and Experimental Setup
@@
4.1.1 MFNet [4]
4.1.2 PST900 [1]
Table 2: Quantitative comparisons (%) with the SoTA RGB-T scene parsing methods on the PST900 test set. The symbol “-” denotes missing data in the original publication, and the best results are presented in bold font.
Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU
MFNet [4] - 98.6 - 41.1 - 64.2 - 60.3 - 20.7 - 57.0
RTFNet [3] - 98.9 - 36.4 - 75.3 - 52.0 - 25.3 - 57.6
EGFNet [60] - 99.2 - 74.3 - 83.0 - 71.2 - 64.6 - 78.5
ABMDRNet [5] - 98.7 - 24.1 - 72.9 - 54.9 - 57.6 - 67.3
FEANet [25] - - - - - - - - - - 91.4 85.5
DBCNet [46] - 98.9 - 62.3 - 71.1 - 52.4 - 40.6 - 74.5
CAINet [10] 99.6 99.5 95.8 80.3 96.0 88.0 88.3 77.2 91.3 78.6 94.2 84.7
EAEFNet [2] 99.8 99.5 92.2 80.4 91.0 87.7 93.0 83.9 79.3 75.6 91.1 85.4
GMNet [9] 99.8 99.4 90.2 85.1 89.0 83.8 88.2 73.7 80.8 78.3 89.6 84.1
CRM-RGBTSeg [7] - 99.6 - 79.5 - 89.6 - 89.0 - 82.2 - 88.0
Table 3: Quantitative comparisons (%) with the SoTA RGB-T scene parsing methods on the KP Day-Night test set. The best results are presented in bold font. Although results for Acc and IoU for certain categories are omitted in the table, they are still included in the calculations of the corresponding mean values.
@@ Methods
Table 4: Quantitative comparisons (%) with the SoTA data-fusion scene parsing networks on the NYU-Depth V2 test set. The symbol “-” denotes missing data in the original publication, and the best results are presented in bold font.
@@
Table 5: An ablation study (%) on different data input strategies on the MFNet test set.
@@
Table 6: An ablation study (%) on different CSPD construction blocks on the MFNet test set.
@@
Duplex ResNet-101 [14] sharing weights 66.3 48.0 38.4 19.9 9.1 0.0 29.6 23.6 42.3 36.8
Duplex Swin-Transformer-S [15] sharing weights 89.7 74.1 67.2 47.6 31.0 6.3 52.7 57.0 74.6 58.2
Duplex MiT-B4 sharing weights [41] sharing weights 87.8 74.7 64.9 48.4 40.9 13.1 51.4 48.1 73.2 58.6
Duplex ConvNeXt-S [17] with separate weights 89.6 74.6 66.6 46.3 45.7 10.2 56.6 62.5 73.6 61.2
Duplex ConvNeXt-S [17] sharing weights (Ours) 89.1 76.0 67.2 50.6 41.3 7.9 58.6 58.7 75.1 60.9
Table 7: An ablation study (%) on the effectiveness of GLCA and CCG on the MFNet test set. When both components are removed, the spatial priors extracted using CSPD and the global context extracted using ViT are fused via element-wise summation after resolution alignment.
@@
Table 8: An ablation study (%) on different symmetric and asymmetric encoder architectures on the MFNet test set.
Methods (encoder) Car Person Bike Curve Car Stop Guardrail Color Cone Bump mAcc mIoU
@@
Duplex BEiTv2-B/16 [21] concatenation 65.7 57.4 12.7 27.7 8.5 0.0 14.7 20.6 40.9 33.7
Duplex BEiTv2-B/16 [21] summation 65.4 57.1 23.3 28.0 8.5 0.0 19.9 16.2 42.6 35.0
Duplex ConvNeXt-S [17] concatenation 89.7 74.1 66.4 45.3 29.8 4.8 52.8 55.5 69.2 57.4
Duplex ConvNeXt-S [17] summation 90.7 74.6 66.9 46.8 32.9 6.6 56.6 52.7 72.0 58.5
HAPNet CMNeXt-B4 [48] 88.5 74.0 62.5 44.9 41.1 12.0 53.3 59.0 71.6 59.3
@@
Table 9: An ablation study (%) on different VFMs on the MFNet test set. “MM” represents multi-modal pre-training. BEiT and BEiTv2 are trained via a self-supervised learning strategy named masked image modeling, while DINOv2 is trained via discriminative self-supervised learning strategy.
@@
DeiT-B/16 [69] Supervised ImageNet-1K 90.4 75.2 66.6 47.6 36.3 0.7 59.9 60.0 68.3 59.4
AugReg-B/16 [70] Supervised ImageNet-22K 90.3 74.9 66.8 48.2 35.8 6.0 56.6 56.5 74.7 59.3
BEiT-B/16 [20] Self-Supervised ImageNet-22K 90.7 75.0 67.6 48.6 39.1 4.2 55.0 59.0 72.5 59.7
DINOv2-B/14 [22] Self-Supervised LVD-142M (MM) [22] 90.0 74.7 66.8 45.1 42.1 9.7 56.2 59.7 76.4 60.3
BEiTv2-B/16 [21] Self-Supervised ImageNet-22K (MM) 89.1 76.0 67.2 50.6 41.3 7.9 58.6 58.7 75.1 60.9
@@ 4.1.4 NYU-Depth V2 [34]
This dataset has been widely used for the evaluation of indoor scene parsing performance. It provides $1,449$ pairs of RGB and depth images (resolution: $480\times 640$ pixels), and semantic annotations for $40$ categories. We adopt the same dataset splitting strategy as presented in [51].
4.2 Evaluation Metrics